OpenStudy (anonymous):

let f(x)=8x^3-22x^2-4 and g(x)=4x-3. Find f(x)/g(x)

OpenStudy (bahrom7893):

\[f(x)=8x^3-22x^2-4\]\[g(x)=4x-3\]\[ \frac{f(x)}{g(x)}-?\]

OpenStudy (bahrom7893):

\[\frac{f(x)}{g(x)}=\frac{8x^3-22x^2-4}{4x-3}\] factor the top and tell me what you get.